The market value balance sheet for Bobaflex Manufacturing is shown here. The company has declared a 20 percent stock dividend. The stock goes ex dividend tomorrow (the chronology for a stock dividend is similar to that for a cash dividend).

  

Market Value Balance Sheet
  Cash $ 92,000   Debt $ 151,000
  Fixed assets 760,000   Equity 701,000
     Total $ 852,000      Total $ 852,000

  

There are 27,000 shares of stock outstanding. What is the current share price? (Do not round intermediate calculations. Round your answer to 2 decimal places, e.g., 32.16.)

  

  Current stock price $   

  

What will the ex-dividend price be? (Do not round intermediate calculations. Round your answer to 2 decimal places, e.g., 32.16.)

   Ex-dividend price $   
Answer #1

a. Current Share price = Total equity/ Number of shares = 701,000/27,000 = $25.96

b. Since the stock gives 20% stock dividend, the ex-dividend stock price = 25.96*(1-0.2) = $20.77